Monica is four years older than Kate. Five years ago, she was twice as old as Kate was. Find their ages now.
Answer by macston(5194)

K=Kate's age; M=Monica's age=K+4yrs
M-5yrs=2(K-5yrs) Substitute for M.
K+4yrs-5yrs=2K-10yrs Subtract K from each side.
-1yrs=K-10 yrs Add 10yrs to each side.
9yrs=K ANSWER 1: Kate is 9 years old.
M=K+4yrs=13yrs+4yrs=13yrs ANSWER 2: Monica is 13 years old
CHECK:
M-5yrs=2(K-5yrs)
13yrs-5yrs=2(9yrs-5yrs)
8yrs=2(4yrs)
8yrs=8yrs